Space Stations

NASA International Space Station On-Orbit Status 5 November 2015

By Marc Boucher
Status Report
November 10, 2015
Filed under , ,
NASA International Space Station On-Orbit Status 5 November 2015
NASA Astronauts Scott Kelly (left) and Kjell Lindgren will conduct a spacewalk today to upgrade and service the International Space Station. Credit: NASA.

Two astronauts are ready to finish a cooling system repair job that was started three years ago on the outside of the International Space Station. NASA astronauts Scott Kelly and Kjell Lindgren will exit the U.S. Quest airlock Friday at 7:10 a.m. EST for about six hours and 30 minutes of spacewalking tasks.
Kelly and Lindgren are restoring the port truss cooling system back to its original configuration after leak repair work done in November 2012 by spacewalkers Suni Williams and Aki Hoshide. More leak repair work was done to the system during a May 2013 spacewalk by NASA astronauts Chris Cassidy and Tom Marshburn.

Kelly and Lindgren will also top off the cooling system’s ammonia levels using storage tanks. The Photovoltaic Thermal Control System dissipates heat generated at the space station from radiators attached to the truss structure. This will be the second spacewalk for both astronauts whose first was on Oct. 28.

Japanese astronaut Kimiya Yui and veteran cosmonaut Sergey Volkov will assist the spacewalkers into their spacesuits and the airlock before depressurization begins. Yui and Volkov will also welcome the spacewalkers back into the station at the end of their excursion.

On-Orbit Status Report

Crew Reduced Duty Day: Today was a reduced duty day for the US crew in advance of tomorrow’s EVA.

Extravehicular Activity (EVA) Preparations: The US crew members conducted final preparations to the Equipment Lock, Extravehicular Mobility Units (EMUs), and ancillary hardware. They also reviewed related procedures and briefing packages then conducted a conference with the ground to answer any questions. US EVA #33 (P6 Return To Original Configuration) is scheduled to occur tomorrow with hatch opening taking place at approximately 6:15am CST.

Reaction Self Test: Kelly completed two sessions today for the Reaction Self Test investigation, one after he woke up in the morning and one prior to his sleep period. These sessions, and an additional late session on Friday, were scheduled to provide data in close proximity to the extravehicular activity (EVA) planned for this Friday. Reaction Self Test helps crewmembers objectively identify when their performance capability is degraded by various fatigue-related conditions that can occur as a result of ISS operations and time in space (e.g., acute and chronic sleep restriction, slam shifts, EVA, and residual sedation from sleep medications).

Space Headaches: Yui and Lindgren completed their weekly Space Headaches questionnaires today. Headaches can be a common complaint during spaceflight. The Space Headaches experiment will provide information that may help in the development of methods to alleviate associated symptoms and improvement in the well-being and performance of crew members in space.

Urine Processing Assembly (UPA) Status: Yesterday, the crew transferred a tank (EDV) of Russian pretreated urine to the UPA Wastewater Storage Tank Assembly (WSTA). The quantity transferred was unexpected. Today, the crew reported that when they tried to perform another EDV transfer, the transfer hose had air bubbles in it and liquid was not flowing. The problem appears to be a leak in the bladder of the EDV that was transferred yesterday which allowed air to enter the hose, the WSTA and the recycle tank. To remove as much of the air as possible from the UPA, the crew drained the WSTA and the recycle tank that was removed yesterday as well as the recycle tank that had been installed today. They also replaced the EDV transfer hose.

Food Warmer Inspection: Yui performed an annual inspection of the US Food Warmer located within the Service Module. While performing this activity, he photo documented wear indications and succesfully performed a powered functional test.

Today’s Planned Activities
All activities were completed unless otherwise noted.

Reaction Self-Test (morning)
Extravehicular Activity (EVA) Reminder for EVA In-Suit Light Exercise (ISLE) Preparation
ISS crew and ???? (RSA Flight Control Management Team) weekly conference
COSMOCARD. Setup. Starting 24-hr ECG Recording
Repairs of SM Interior Panel 317 by Installing Overlay Panels – Day 1
Pre-EVA Crew Periodic Health Status Examination
Filling (separation) of EDV (KOV) for Elektron or EDV-SV. Inspect and separate EDV-SV No.1002 ?.37 (00053403R, ???1???_1_107). For separation use EDV (KOV) No.1199 (00065916R, ???1???_1_109)
Crew Health Status before USOS EVA – Data Transfer
??? Maintenance. Flushing ???
ECLSS Recycle Tank Remove and Replace
On MCC Go ???? Rate Sensor (???) control sub-network user comm check with RS3, RS2, RS1 Laptops.
Ventilation grille cleaning on FGB interior panels (panels 201, 301, 401)
HAM radio session from Columbus [Deferred]
WRS – Recycle Tank Fill
USOS EVA Procedure Review
Final printout of USOS EVA procedures
WRS – Recycle Tank Fill
Pille Sensors prep for USOS EVA
TV Conference with the Participants of Voskhod camp space tour
PROBOY. RS?1 Laptop Ops
Rubber valve replacement in dining table on ISS RS
PROBOY. Penetration Simulator Ops
IMS and Stowage Conference
Removal of depress hose for nominal operations [Deferred]
EVA Equipment Lock Preparation
EXPOSE-R. Copy and Downlink Data
WRS – Recycle Tank Fill
Cleaning FGB ??1 Circulation Fan Screens
USOS EVA Procedure Conference
PROBOY. Copy and Downlink Data
CONTENT. Experiment Ops
Cleaning fan screens on FGB interior panels 116, 316, 231, 431
SHD – Questionnaire
Private medical conference before EVA from USOS
Post-Exercise Hygiene Procedure
CONTENT. Experiment Ops
SHD – Questionnaire
Reaction Self-Test (evening)
Preparation of Reports for Roscosmos Web Site and Social Media
ECON-M. Observation and Photography
URAGAN Observations and Photography
IMS Delta File Prep

Completed Task List Items

Ground Activities
All activities were completed unless otherwise noted.
ETCS Loop B Preparation in support of P6 RTOC EVA

Three-Day Look Ahead:
Friday, 11/06: P6 RTOC EVA
Saturday, 11/07: Post EVA Activities
Sunday, 11/08: Weekly Cleaning, Crew Off Duty

QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – Off
Vozdukh – Manual
[???] 1 – SM Air Conditioner System (“SKV1”) – On
[???] 2 – SM Air Conditioner System (“SKV2”) – Off
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Shutdown
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.